OpenStudy (anonymous):

find the first derivative for y=3/(cosx) + 2/(tanx)

OpenStudy (shivaniits):

\[y=3\sec x+2\cot x\] \[dy/dx=(3\tan x \times \sec x)-2(\csc x)^{2}\]

OpenStudy (anonymous):

What is the rule to figure this out. Because I thought that the derivative of cosx is -sinx and the derivative of tanx is sec^2x Do you do the quotient rule for both numbers first and then do the sum rules? : 0(cosx)-3(-sinx) = 3sinx 0(tanx)-2(sec^2x) = -2sec^2x Am I doing it right so far?

OpenStudy (anonymous):

Oh nevermind I just saw that the derivative of 1/cosx is tanxsecx and the derivative of 1/tanx is -csc^2x Thank you so much for your help

OpenStudy (shivaniits):

how to figure out the derivative of cot x \[f(x)=\cot x=\cos x/\sin x\] \[f'(x)=[sinx (-sinx)-cosx(cosx)]/(sinx)^2\] \[f'(x)=[-\sin^2x-\cos^2x]/(sinx)^2\] \[f'(x)=-(\sin^2x +\cos^2x)/\sin^2x\] \[f'(x)=-(1/\sin^2 x)\] \[f'(x)=-\csc^2x \] another solution \[f(x)=\cot x=1/tanx \] \[f'(x)=[tanx(0)-1(\sec^2x)]/\tan^2x \] \[f'(x)=-\sec^2x/\tan^2x \] \[f'(x)=-(1/\cos^2x)(\cos^2x/\sin^2x) \] \[f'(x)=-1/\sin^2x=-\csc^2x \] how to figure out the derivative for sec x \[secx=1/cosx \] \[(secx)'=[1'cosx-1.cosx']/cosx² \] \[(secx)'=[0.cosx-1.(-sinx)]/cosx² \] \[(secx)'=[sinx]/cosx² \] \[(secx)'=[sinx]/cosx].1/cosx \] \[(secx)'= \tan x \times \sec x\]